OpenStudy (anonymous):

\[5- a-1/a+2 = a^2 -1/a+2\] solve

OpenStudy (anonymous):

a^2 +a=5 ?

OpenStudy (anonymous):

i don't believe so.

OpenStudy (anonymous):

a = ?

OpenStudy (anonymous):

a=-2.79 ?

OpenStudy (anonymous):

a^2 + a = 5 is correct a^2 + a - 5 = 0 this can be solved using formula a = 1.79, -2.79

OpenStudy (anonymous):

that is incorrect as well.

OpenStudy (anonymous):

I really have problems with these ones. The correct answer is 6

OpenStudy (anonymous):

Thank you all for your efforts. I will be posting a similar question in a while.

OpenStudy (anonymous):

the answer is not 6 - try plugging 6 into the equation

OpenStudy (anonymous):

the assignment gave me the answer 6 once I got it incorrect 3 times